# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ 5.1. Sturmian sequences
The slope $\alpha$ irrational rotation is the Kronecker system $(\mathbb{T},\mathbb{Z},R_{\alpha})$ defined $R_{\alpha}u:=u+\alpha\!\!\mod 1$. It satisfies $\Sigma_{R_{\alpha}}=\mathbb{Z}[\alpha]/\mathbb{Z}\subseteq\mathbb{T}\sim\widehat{\mathbb{Z}}$, where $\mathbb{Z}[\alpha]$ is the ring generated by $\alpha$, and $\mathbb{Z}[\alpha]/\mathbb{Z}$ is its projection to the circle. A Sturmian sequence $z=(z_{n})_{n\in\mathbb{Z}}\in\{0,1\}^{\mathbb{Z}}$ is a coding of $u\in\mathbb{T}$ by $R_{\alpha}$. It is defined $z_{n}=z(u)_{n}:=\chi_{I}(R^{n}_{\alpha}u)$, where $I$ is either $I^{+}:=[1-\alpha,1)$ or $I^{-}:=(1-\alpha,1]$. We let $Z=\{(z(u)_{n})_{n\in\mathbb{Z}}:u\in\mathbb{T}\}$ with $\phi:Z\to\mathbb{T}$ the factor map satisfying $\phi(z(u))=u$. The following lemma summarizes some well known facts.
@@
The $\mathbb{Z}$ Sturmian subshift $(Z,\mathbb{Z},S)$ is strictly ergodic almost automorphic with $\Sigma_{S}=\mathbb{Z}[\alpha]/\mathbb{Z}$, $\mathfrak{M}_{T}=\{1,2\}$, and $\mathbb{T}_{1}:=\mathbb{T}\backslash\{R^{n}_{\alpha}0:n\in\mathbb{Z}\}$.
@@ 5.2. Grid patches
Let ${{\boldsymbol{f}}}_{0}:=(0,\sec(2\pi/20))$ and ${{\boldsymbol{f}}}_{1}:=(1,-\tan(2\pi/20))$, and let $A=[{\boldsymbol{f}}_{0}^{t};{\boldsymbol{f}}_{1}^{t}]$ be the matrix with ${\boldsymbol{f}}_{0}$ and ${\boldsymbol{f}}_{1}$ as rows. Define a rhombus
@@
For any ${\boldsymbol{u}}\in\mathbb{T}^{5}_{0}$, the 00 and $1$-grid lines in $y({\boldsymbol{u}})$ partition ${\mathbb{R}}^{2}$ into a tessellation $\{R_{\boldsymbol{n}}\}_{{\boldsymbol{n}}\in\mathbb{R}^{2}}$ by congruent rhombs. Here,
@@
where ${\boldsymbol{u}}^{\prime}=(u_{0},u_{1})\in[0,1)^{2}$. Note that ${{\boldsymbol{0}}}\in R_{{\boldsymbol{0}}}$. We define the grid patch $g_{{\boldsymbol{n}}}({{\boldsymbol{u}}})$ to be the patch in $y({{\boldsymbol{u}}})$ supported on the rhomb $R_{{\boldsymbol{n}}}$ (see Figure 4 (b)).
@@
Let $y({\boldsymbol{u}})\in Y_{1}$ (so all crossings are $2$-fold). For ${{\boldsymbol{n}}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, define the symbol of $g_{{\boldsymbol{n}}}({{\boldsymbol{u}}})$, denoted $\sigma(g_{{\boldsymbol{n}}}({{\boldsymbol{u}}}))=z_{n_{0}}\otimes z^{\prime}_{n_{1}}\in\{0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1\}$ as follows: If a single $4$-grid line passes through $R_{{\boldsymbol{n}}}$, put $z_{n_{0}}=0$, but if two $4$-grid lines pass through $R_{{\boldsymbol{n}}}$, put $z_{n_{0}}=1$. Similarly, if a single $2$-grid line passes through $R_{{\boldsymbol{n}}}$, put $z^{\prime}_{n_{1}}=0$, but if two $2$-grid lines pass through $R_{{\boldsymbol{n}}}$, put $z^{\prime}_{n_{1}}=1$.
@@
Suppose ${{\boldsymbol{u}}}\in\mathbb{T}^{5}_{0,1}$ satisfies $u_{0}=u_{1}=0$. Then $z_{0}\otimes z^{\prime}_{0}=\sigma(g_{{\boldsymbol{0}}}({{\boldsymbol{u}}}))$ is given by $z_{0}=\chi_{[1-\alpha,1)}(u_{4})$ and $z^{\prime}_{0}=\chi_{[1-\alpha,1)}(u_{2})$. Moreover, $\{\sigma(g_{{\boldsymbol{n}}}({{\boldsymbol{u}}}))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}=z(u_{4})\otimes z(u_{2})$.
@@
Figure 8. Left: In the case $z^{\prime}_{0}=0$, a single 2-grid line, $\ell_{2,0}(u_{2})$, crosses $R_{{\boldsymbol{0}}}$, so that $u_{2}\in[0,1-\alpha)$ since $\ell_{2,0}(u_{2})$ is left of $\ell_{2,0}(1-\alpha)$. Right: In the case $z^{\prime}_{0}=1$, two 2-grid lines cross $R_{{\boldsymbol{0}}}$ and $u_{2}\in[1-\alpha,1)$.
@@ 5.3. Duals of grid patches
Suppose $x\in\mathsf{X}_{1}$ ((((a nonsingular Penrose tiling)))) with $x=y^{*}$ for $y=y({\boldsymbol{u}})\in Y_{1}$. Suppose ${\boldsymbol{b}}_{\boldsymbol{n}}\in v(y)$ satisfies ${\boldsymbol{b}}_{\boldsymbol{n}}=\ell_{0,n_{0}}\cap\ell_{1,n_{1}}$, ${\boldsymbol{n}}=(n_{0},n_{1})$, so ${\boldsymbol{b}}_{\boldsymbol{n}}$ is the lower left vertex of the rhombus $R_{\boldsymbol{n}}$ in (5.2). Let $E_{\boldsymbol{n}}\in y$ be the tile in $g_{\boldsymbol{n}}({\boldsymbol{u}})$ with vertex ${\boldsymbol{b}}_{\boldsymbol{n}}$. The dual $D_{\boldsymbol{n}}={\boldsymbol{b}}_{\boldsymbol{n}}^{*}\in x$ of ${\boldsymbol{b}}_{\boldsymbol{n}}$ is a type ${\boldsymbol{v}}^{\prime}_{0}\wedge{\boldsymbol{v}}^{\prime}_{1}$ tile from $p_{5}$ in $x$, and the dual ${\boldsymbol{d}}_{\boldsymbol{n}}=E_{\boldsymbol{n}}^{*}$ of the tile $E_{\boldsymbol{n}}\in y$ is the upper right vertex of $D_{\boldsymbol{n}}$. Then $||{\boldsymbol{b}}_{\boldsymbol{n}}-{\boldsymbol{d}}_{\boldsymbol{n}}||<(1/5)(1+\sqrt{5})$.
@@
Use Lemma 4.5 to translate ${\boldsymbol{b}}_{\boldsymbol{n}}$ to the origin, and making it ${\boldsymbol{b}}_{\bf 0}$, by replacing ${\boldsymbol{y}}$ with $K^{-{\boldsymbol{b}}_{\boldsymbol{n}}}{\boldsymbol{u}}$. Then $\{{\boldsymbol{u}}\}=(0,0,u_{2},\{-(u_{2}+u_{4})\},u_{4})$, and we show $||{\boldsymbol{d}}_{\bf 0}||<(1/2)(1+\sqrt{5})$. By (4.3), ${\boldsymbol{d}}_{\bf 0}=E^{*}_{\bf 0}:=(2/5)W^{t}({\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})+\{{\boldsymbol{u}}\})$ for any ${\boldsymbol{s}}\in{\rm int}(E_{\bf 0})$. Making ${\boldsymbol{s}}>0$ small, we see from (4.2) that ${\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})={{\boldsymbol{0}}}$ for ${\boldsymbol{s}}\in{\rm int}(E_{\bf 0})$, since ${\boldsymbol{m}}$ is constant on tiles. Thus $||{\boldsymbol{d}}_{\bf 0}||=(2/5)||W^{t}\{{\boldsymbol{u}}\}||\leq(2/5)||{\boldsymbol{v}}_{2}+{\boldsymbol{v}}_{3}+{\boldsymbol{v}}_{4}||=(2/5)(1+\sqrt{5})/2=(1/5)(1+\sqrt{5})$. ∎
For $x\in X_{p_{5}}$ (in particular, for $x\in\mathsf{X}$), we call an infinite sequence $\{\dots,D_{-1},D_{0},D_{1},\dots\}$ of tiles in $x$ a forward orientation, direction ${\boldsymbol{v}}_{j}^{\perp}$ pseudo-line each pair $D_{i}$ and $D_{i+1}$ are adjacent across a type ${\boldsymbol{v}}^{\prime}_{j}$ edge in the direction $R_{2\pi/4}{\boldsymbol{v}}_{j}$. For $x({\boldsymbol{u}})\in\mathsf{X}_{1}$ nonsingular, every direction ${\boldsymbol{v}}_{j}^{\perp}$ pseudo-line is the dual of the $2$-fold crossings along some grid line $\ell_{j,k}$ in $y({\boldsymbol{u}})$. A pseudo-line segment is a contiguous finite subsequence $\{D_{k},D_{k+1},\dots,D_{k+n-1}\}$ of a pseudo-line.
@@
Fix a Penrose tiling $x\in\mathsf{X}$. Then there is an injective map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}:\mathbb{Z}^{2}\to x$ that is onto the set of all type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}_{1}^{\prime}$ tiles in $D\in x$, such that for each ${\boldsymbol{n}}\in\mathbb{Z}^{2}$, $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$ is forward along a direction ${\boldsymbol{v}}_{0}^{\perp}$ pseudo-line from $D_{\boldsymbol{n}}$, and $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$ is forward along a direction ${\boldsymbol{v}}_{1}^{\perp}$ pseudo-line from $D_{\boldsymbol{n}}$. This map is unique for $x\in\mathsf{X}_{1}$ if we assume $D_{\bf 0}={\boldsymbol{b}}_{\bf 0}^{*}$, or for $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ if we assume $D_{\bf 0}=({\boldsymbol{b}}^{\prime}_{\bf 0})^{*}$ with $x^{\prime}\in\mathsf{X}_{1}$ and $\rho(x,x^{\prime})$ sufficiently small.
@@
If $x\in\mathsf{X}_{1}$ is nonsingular, the map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}$ defined in Lemma 5.5 satisfies the lemma. If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ is singular, there is a sequence $x_{k}\in\mathsf{X}_{1}$ so that $x_{k}\to x$ in the metric $\rho$, (i.e., $\mathsf{X}=\overline{\mathsf{X}_{1}}$, Remark 4.7). Let $x_{k}=y({\boldsymbol{u}}_{k})^{*}$ and let ${\boldsymbol{n}}\mapsto D^{k}_{\boldsymbol{n}}$ be the map defined in Lemma 5.5. It satisfies $D^{k}_{\boldsymbol{n}}={\boldsymbol{b}}^{k}_{\boldsymbol{n}}$ for ${\boldsymbol{b}}^{k}_{\boldsymbol{n}}\in v(y({\boldsymbol{u}}_{k}))$. Let $r_{n}>0$ be an increasing sequence, and for each $n$, choose $K_{n}$ large enough that $k\geq K_{n}$ implies $x_{k}[B_{r_{n}}({\bf 0})]=x[B_{r_{n}}({\bf 0})]$. Define $D_{\boldsymbol{n}}:=D^{k}_{\boldsymbol{n}}$ where $r_{n}>||{\boldsymbol{n}}||$ and $k\geq K_{n}$. ∎
@@
For a Penrose tiling $x\in\mathsf{X}$ and ${\boldsymbol{n}}\in\mathbb{Z}^{2}$, we define a patch $w_{\boldsymbol{n}}(x)$ in $x$, called a Wang patch, as follows: Starting at $D_{\boldsymbol{n}}$, go forward along the ${\boldsymbol{v}}^{\perp}_{0}$ pseudo-line segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$, then go forward along the ${\boldsymbol{v}}^{\perp}_{1}$ pseudo-line segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$, then go backward along the ${\boldsymbol{v}}^{\perp}_{0}$ pseudo-line segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$, and finally go backward along the ${\boldsymbol{v}}^{\perp}_{1}$ pseudo-line segment, back to $D_{\boldsymbol{n}}$. The patch $w_{\boldsymbol{n}}(x)$ is defined to be this closed pseudo-line path, and all the tiles inside it.
The Wang patches $\{w_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ in $x\in\mathsf{X}$ cover $\mathbb{R}^{2}$ in the sense that $x=\bigcup_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}w_{\boldsymbol{n}}(x)$. We call the tiles $D_{\boldsymbol{n}},D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}},D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$ and $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$ the corners of a Wang patch $w_{\boldsymbol{n}}(x)$, and we call the pseudo-line segments from corner to corner, its margins; bottom: ${\boldsymbol{\mu}}_{\boldsymbol{n}}^{\rm b}$, and top: ${\boldsymbol{\mu}}_{\boldsymbol{n}}^{\rm t}$, left: ${\boldsymbol{\mu}}_{\boldsymbol{n}}^{\rm l}$, and right: ${\boldsymbol{\mu}}_{\boldsymbol{n}}^{\rm r}$. Adjacent Wang patches in $x$ are share common margins: $w_{\boldsymbol{n}}(x)\cap w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}(x)={\boldsymbol{\mu}}^{\rm r}_{\boldsymbol{n}}={\boldsymbol{\mu}}^{\rm l}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$ and $w_{\boldsymbol{n}}(x)\cap w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}(x)={\boldsymbol{\mu}}^{\rm t}_{\boldsymbol{n}}={\boldsymbol{\mu}}^{\rm b}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$. The following is a consequence of Lemma 5.6.
@@
For $x\in X_{1}$, let $\varphi(x)=y({\boldsymbol{u}})$ ((((all crossings $2$-fold)))), so that $x=y({\boldsymbol{u}})^{*}$. Then each Wang patch in $x$ is the dual of the corresponding grid patch: $w_{\boldsymbol{n}}(x)=g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$.

### v2

@@ 5.1. Sturmian sequences
Recall that $\alpha:=\gamma-1\approx 0.6180$. The slope $\alpha$ irrational rotation is the Kronecker system $(\mathbb{T},\mathbb{Z},R_{\alpha})$ defined by $R_{\alpha}u:=u+\alpha\!\!\mod 1$. It satisfies $\Sigma_{R_{\alpha}}=\mathbb{Z}[\alpha]/\mathbb{Z}\subseteq\mathbb{T}\sim\widehat{\mathbb{Z}}$, where $\mathbb{Z}[\alpha]$ is the ring generated by $\alpha$, and $\mathbb{Z}[\alpha]/\mathbb{Z}$ is its projection to the circle. A Sturmian sequence $z=(z_{n})_{n\in\mathbb{Z}}\in\{0,1\}^{\mathbb{Z}}$ is a coding of $u\in\mathbb{T}$ by $R_{\alpha}$. It is defined by $z_{n}=z(u)_{n}:=\chi_{I}(R^{n}_{\alpha}u)$, where $I$ is either $I^{+}:=[1-\alpha,1)$ or $I^{-}:=(1-\alpha,1]$. We let $Z=\{(z(u)_{n})_{n\in\mathbb{Z}}:u\in\mathbb{T}\}$, with $\phi:Z\to\mathbb{T}$ being the factor map satisfying $\phi(z(u))=u$. The following lemma summarizes some well known facts.
@@
The $\mathbb{Z}$ Sturmian subshift $(Z,\mathbb{Z},S)$ (not SFT) is strictly ergodic almost automorphic with $\Sigma_{S}=\mathbb{Z}[\alpha]/\mathbb{Z}$, $\mathfrak{M}_{T}=\{1,2\}$, and $\mathbb{T}_{1}:=\mathbb{T}\backslash\{R^{n}_{\alpha}0:n\in\mathbb{Z}\}$.
@@ 5.2. Grid patches
Let
${{\boldsymbol{f}}}_{0}$ $:=(0,\sec(2\pi/20))=(0,1+\sqrt{5})\approx(0,3.236),{\rm\ and}$
${{\boldsymbol{f}}}_{1}$ $:=(1,-\tan(2\pi/20))=(1,\sqrt{5+2\sqrt{5}})\approx(-1,3.078),$ (5.1)
and let $A$ be the matrix with ${\boldsymbol{f}}_{0}$ and ${\boldsymbol{f}}_{1}$ as columns. We will define a rhombic tile $R$ by
@@
[…]
@@
where in the second equality, we assume without loss of generality that ${\boldsymbol{u}}^{\prime}:=(u_{0},u_{1})\in[0,1)^{2}$.
Note that ${{\boldsymbol{0}}}\in R_{{\boldsymbol{0}}}$. We define the grid patch $g_{{\boldsymbol{n}}}({{\boldsymbol{u}}})$ to be the patch in $y({{\boldsymbol{u}}})$ supported on the rhombus $R_{{\boldsymbol{n}}}$ (see Figure 4(b)).
@@
Let $y({\boldsymbol{u}})\in Y_{1}$ be a nonsingular grid tiling (all crossings $2$-fold). For ${{\boldsymbol{n}}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, define the symbol of the grid patch $g_{{\boldsymbol{n}}}({{\boldsymbol{u}}})$, denoted by $\sigma(g_{{\boldsymbol{n}}}({{\boldsymbol{u}}}))=z_{n_{0}}\otimes z^{\prime}_{n_{1}}\in\{0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1\}$ as follows: If a single $4$-grid line passes through $R_{{\boldsymbol{n}}}$, put $z_{n_{0}}=0$, but if two $4$-grid lines pass through $R_{{\boldsymbol{n}}}$, put $z_{n_{0}}=1$. Similarly, if a single $2$-grid line passes through $R_{{\boldsymbol{n}}}$, put $z^{\prime}_{n_{1}}=0$, but if two $2$-grid lines pass through $R_{{\boldsymbol{n}}}$, put $z^{\prime}_{n_{1}}=1$.
@@
Suppose that ${{\boldsymbol{u}}}\in\mathbb{T}^{5}_{0,1}$ satisfies $u_{0}=u_{1}=0$. Then $z_{0}\otimes z^{\prime}_{0}=\sigma(g_{{\boldsymbol{0}}}({{\boldsymbol{u}}}))$ is given by $z_{0}=\chi_{[1-\alpha,1)}(u_{4})$ and $z^{\prime}_{0}=\chi_{[1-\alpha,1)}(u_{2})$. Moreover, $\{\sigma(g_{{\boldsymbol{n}}}({{\boldsymbol{u}}}))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}=z(u_{4})\otimes z(u_{2})$.
@@
Figure 8. (a) In the case $z^{\prime}_{0}=0$, a single 2-grid line, $\ell_{2,0}(u_{2})$, crosses $R_{{\boldsymbol{0}}}$, so that $u_{2}\in[0,1-\alpha)$ since $\ell_{2,0}(u_{2})$ is left of $\ell_{2,0}(1-\alpha)$. (b) In the case $z^{\prime}_{0}=1$, two 2-grid lines cross $R_{{\boldsymbol{0}}}$ and $u_{2}\in[1-\alpha,1)$. (c) The singular case ($u_{2}=0$).
@@ 5.3. Duals of grid patches
Let $x\in\mathsf{X}_{1}$ be a nonsingular Penrose tiling with $x=y({\boldsymbol{u}})^{*}$ for $y=y({\boldsymbol{u}})\in Y_{1}$. Suppose that ${\boldsymbol{b}}_{\boldsymbol{n}}\in v(y)$ satisfies ${\boldsymbol{b}}_{\boldsymbol{n}}=\ell_{0,n_{0}}\cap\ell_{1,n_{1}}$, ${\boldsymbol{n}}=(n_{0},n_{1})$, so ${\boldsymbol{b}}_{\boldsymbol{n}}$ is the lower left vertex of the rhombus $R_{\boldsymbol{n}}$ in (5.3). Let $E_{\boldsymbol{n}}\in y$ be the tile in $g_{\boldsymbol{n}}({\boldsymbol{u}})$ with vertex ${\boldsymbol{b}}_{\boldsymbol{n}}$. The dual $D_{\boldsymbol{n}}={\boldsymbol{b}}_{\boldsymbol{n}}^{*}\in x$ of ${\boldsymbol{b}}_{\boldsymbol{n}}$ is a type ${\boldsymbol{v}}^{\prime}_{0}\wedge{\boldsymbol{v}}^{\prime}_{1}$ tile from $p_{5}$ in $x$, and the dual ${\boldsymbol{d}}_{\boldsymbol{n}}=E_{\boldsymbol{n}}^{*}$ of the tile $E_{\boldsymbol{n}}\in y$ is the upper right vertex of $D_{\boldsymbol{n}}$. Then $||{\boldsymbol{b}}_{\boldsymbol{n}}-{\boldsymbol{d}}_{\boldsymbol{n}}||<(1/5)(1+\sqrt{5})\approx 0.647$.
@@
Use Lemma 4.6 to translate ${\boldsymbol{b}}_{\boldsymbol{n}}$ to the origin, and making it ${\boldsymbol{b}}_{\bf 0}$, by replacing ${\boldsymbol{u}}$ with $K^{-{\boldsymbol{b}}_{\boldsymbol{n}}}{\boldsymbol{u}}$. Then $\{{\boldsymbol{u}}\}=(0,0,u_{2},\{-(u_{2}+u_{4})\},u_{4})$ with $u_{2},u_{4}\in[0,1)$. We show $||{\boldsymbol{d}}_{\bf 0}||<(1/2)(1+\sqrt{5})$. By (4.3), ${\boldsymbol{d}}_{\bf 0}=E^{*}_{\bf 0}:=(2/5)W^{t}({\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})+\{{\boldsymbol{u}}\})$ for any ${\boldsymbol{s}}\in{\rm int}(E_{\bf 0})$. Making ${\boldsymbol{s}}>0$ small, we see from (4.2) that ${\boldsymbol{m}}({\boldsymbol{s}},{\boldsymbol{u}})={{\boldsymbol{0}}}$ for ${\boldsymbol{s}}\in{\rm int}(E_{\bf 0})$, since ${\boldsymbol{m}}$ is constant on tiles. Thus $||{\boldsymbol{d}}_{\bf 0}||=(2/5)||W^{t}\{{\boldsymbol{u}}\}||\leq(2/5)||{\boldsymbol{v}}_{2}+{\boldsymbol{v}}_{3}+{\boldsymbol{v}}_{4}||=(2/5)(1+\sqrt{5})/2=(1/5)(1+\sqrt{5})$. ∎
Let $x\in\mathsf{X}$ (or more generally, $x\in X_{p_{5}}$). A ${\boldsymbol{v}}_{j}^{\perp}$-trail in $x$ is an infinite sequence of tiles $\{\dots,D_{-1},D_{0},D_{1},\dots\}$ in $x$ such that each $D_{i+1}$ is adjacent across to $D_{i}$ across a type ${\boldsymbol{v}}^{\prime}_{j}$ edge in the direction $R_{2\pi/4}{\boldsymbol{v}}_{j}$. For $x=y({\boldsymbol{u}})^{*}\in\mathsf{X}_{1}$ every ${\boldsymbol{v}}_{j}^{\perp}$-trail is the dual of the sequence of $2$-fold crossings along a grid line $\ell_{j,k}(u_{k})$ in $y({\boldsymbol{u}})$ for some $k\in\mathbb{Z}$. A ${\boldsymbol{v}}_{j}^{\perp}$-trail segment is a contiguous finite subsequence $\{D_{k},D_{k+1},\dots,D_{k+n-1}\}$ of a ${\boldsymbol{v}}_{j}^{\perp}$-trail.
@@
Fix a Penrose tiling $x\in\mathsf{X}$ and let $x^{\prime}$ be the set of all type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}_{1}^{\prime}$ tiles in $D\in x$. Then there is a bijective map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}:\mathbb{Z}^{2}\to x^{\prime}$ such that for each ${\boldsymbol{n}}\in\mathbb{Z}^{2}$, a ${\boldsymbol{v}}_{0}^{\perp}$-trail segment connects $D_{\boldsymbol{n}}$ to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$ and a ${\boldsymbol{v}}_{1}^{\perp}$-trail segment connects $D_{\boldsymbol{n}}$ to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$. The map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}$ is unique for $x\in\mathsf{X}_{1}$ if we assume $D_{\bf 0}={\boldsymbol{b}}_{\bf 0}^{*}$, or for $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ if we assume $D_{\bf 0}=({\boldsymbol{b}}^{\prime}_{\bf 0})^{*}$ with $x^{\prime}\in\mathsf{X}_{1}$ and $\rho(x,x^{\prime})$ sufficiently small.
@@
If $x\in\mathsf{X}_{1}$ is nonsingular, the map ${\boldsymbol{n}}\mapsto D_{\boldsymbol{n}}$ defined in Lemma 5.5 satisfies the lemma. If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ is singular, there is a sequence $x_{k}\in\mathsf{X}_{1}$ so that $x_{k}\to x$ in the metric $\rho$, (i.e., $\mathsf{X}=\overline{\mathsf{X}_{1}}$, Lemma 4.8). Let $x_{k}=y({\boldsymbol{u}}_{k})^{*}$ and let ${\boldsymbol{n}}\mapsto D^{k}_{\boldsymbol{n}}$ be the map defined in Lemma 5.5. It satisfies $D^{k}_{\boldsymbol{n}}={\boldsymbol{b}}^{k}_{\boldsymbol{n}}$ for ${\boldsymbol{b}}^{k}_{\boldsymbol{n}}\in v(y({\boldsymbol{u}}_{k}))$. Let $r_{n}>0$ be an increasing sequence, and for each $n$, choose $K_{n}$ large enough that $k\geq K_{n}$ implies $x_{k}[B_{r_{n}}({\bf 0})]=x[B_{r_{n}}({\bf 0})]$. Define $D_{\boldsymbol{n}}:=D^{k}_{\boldsymbol{n}}$ where $r_{n}>||{\boldsymbol{n}}||$ and $k\geq K_{n}$. ∎
@@
For a Penrose tiling $x\in\mathsf{X}$ and ${\boldsymbol{n}}\in\mathbb{Z}^{2}$, we define a patch $w_{\boldsymbol{n}}(x)$ in $x$, called a Wang patch, as follows: Starting at $D_{\boldsymbol{n}}$, go forward along the ${\boldsymbol{v}}^{\perp}_{0}$-trail segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$ (Lemma 5.6), then go forward along the ${\boldsymbol{v}}^{\perp}_{1}$-trail segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$, then go backward along the ${\boldsymbol{v}}^{\perp}_{0}$-trail segment to $D_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$, and finally go backward along the ${\boldsymbol{v}}^{\perp}_{1}$-trail segment back to $D_{\boldsymbol{n}}$. The patch $w_{\boldsymbol{n}}(x)$ is defined to be this loop of closed trail segments, together with all the tiles inside it.
The Wang patches $\{w_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ in $x\in\mathsf{X}$ cover $\mathbb{R}^{2}$ in the sense that $x=\bigcup_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}w_{\boldsymbol{n}}(x)$. We call the four type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}_{1}^{\prime}$ tiles in $w_{\boldsymbol{n}}(x)$ its corners. The following is a consequence of Lemma 5.6.
@@
For $x\in\mathsf{X}_{1}$, let $x=y({\boldsymbol{u}})^{*}$ where $y({\boldsymbol{u}})\in Y_{1}$. Then each Wang patch in $x$ is the dual of the corresponding grid patch: $w_{\boldsymbol{n}}(x)=g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$.